# CATP-LLM: Empowering Large Language Models for Cost-Aware Tool Planning

### v2

@@ 3.2 Cost-Aware Offline Reinforcement Learning
With the proposed TPL, the LLM is capable to devise various tool plans by simply predicting tool and dependency tokens. To optimize the plan generation process, CATP-LLM then fine-tunes the LLM through the efficient CAORL mechanism. Specifically, CAORL includes: i) a context augmentation scheme to effectively integrate tool cost information into the LLM’s input context; ii) an offline RL-based fine-tuning algorithm that guides the LLM to achieve the optimal performance-cost trade-off in tool planning.
@@ 3.2.1 Context Augmentation
Figure 4: Illustrations of the prompt template and context augmentation design.
The details of the context augmentation scheme are elaborated as follows. First, we categorize the input sizes into $\{1,\cdots,k\}$ levels as two very close sizes will not make a big difference (e.g. resolution 480$\times$520 vs. 490$\times$510, batch size 32 vs. 36). Subsequently, each tool is assigned with a cost attribute vector $c(t_{i})=(c_{1}^{i},\cdots,c_{k}^{i})$, indicating its execution costs at different levels of input sizes. Afterward, we employ a linear layer to extract features from the tool cost attributes. We then encode the importance of cost features by adding the following importance vector:
@@
The remaining problem here is how to acquire the cost attributes of each tool $c(t_{i})$. To tackle this issue, we manually construct various tool plans and ensure that each tool is used at least once in these plans. We then simulate diverse input data at varying size levels, feed the simulated data to each plan, and record the execution costs of each tool (e.g. execution time and memory consumption), which ultimately yields the cost attributes of each tool.
@@ 3.2.2 Learning Through Offline RL
Despite augmenting the LLM with rich input context, achieving cost-aware tool planning remains challenging due to the inherent trade-off between plan performance and costs. Hence, it is imperative to fine-tune the LLM to acquire the domain knowledge for cost-aware planning. Notably, we observe that the selection of a tool can have cascading effects on the subsequent selection. For instance, selecting a powerful yet costly tool may lead to avoiding the selection of expensive tools later to prevent the costs from becoming unaffordable. This suggests that the cost-aware tool planning is essentially a sequential decision problem, a domain where reinforcement learning (RL) [43] excels. Thus, we design the fine-tuning algorithm based on RL, as illustrated in Figure 5 and detailed below.
@@
The primary challenge of implementing such a learning algorithm lies in the collection of training datasets. In lack of public tool plan datasets for fine-tuning, we propose a dataset generation method to collect the necessary training data. This approach involves sampling tasks from a task pool and instructing the LLM (specifically GPT-4 [1] in our practical implementation) to create tool plans to address these tasks. Note that not all generated plans are executable, particularly non-sequential ones due to their complex structures. Hence, we further filter out those invalid plans and utilize the rest as the plan dataset for fine-tuning.
4 OpenCATP Platform
In this section, we describe the OpenCATP platform designed to evaluate the efficacy of LLMs in cost-aware tool planning. OpenCATP is built upon OpenAGI [10], which offers a rich toolkit (e.g. image deblurring), diverse tasks for evaluation (e.g. image editing), as well as objective metrics for performance measurement (e.g. ViT Score [10] and BERT Score [53]). This allows us to focus on cost measurement and evaluation in building OpenCATP.
The key feature of OpenCATP is that it defines a cost model to comprehensively reflect the costs of tool plans, inspired by the Function as a Service (FaaS) [36, 20] platforms. To be specific, FaaS providers have developed a mature pricing model in order to charge for the service of running a function based on the execution time and resource consumption. Hence, we take cues from the FaaS platforms and design a pricing model in OpenCATP to quantify the overall costs of a tool plan, which jointly considers execution time and instant/constant memory consumption (detailed in Appendix §$\S$§B). Based on this model, we introduce a unified metric Quality of Plan (QoP) in OpenCATP to quantify the quality of a tool plan:
@@ 4 OpenCATP Platform
where $P_{task}(p)$ quantifies the task performance of a plan with the OpenAGI APIs, $C_{price}(p)$ calculates the execution prices of a tool plan reflecting its overall costs, and $\alpha\in(0,1)$ is the weight parameter. We employ min-max normalization on these two metrics when calculating QoP to make sure their values fall in the same scale. The min/max values are derived from empirical measurement.
Finally, we introduce more challenging tasks that require non-sequential plans for task solving in OpenCATP. Following the back-instruct method in [38], we construct diverse non-sequential tasks by following steps: (1) create a tool graph based on the toolkit and dependencies between tools; (2) extract a sub-graph representing a non-sequential plan; (3) prompt the GPT-4 to generate the descriptions of a task that can be solved by this plan.
Metrics GPT-3.5 GPT-4 Llama2-7B
Zero-Shot Few-Shot HuggingGPT Zero-Shot Few-Shot HuggingGPT IFT RLTF CATP-LLM
Sequential Planning
Task Performance $\uparrow$ 0.6758 0.6988 0.6645 0.6856 0.6982 0.6661 0.6482 0.6938 0.6611
 Costs $\downarrow$ Exec. Prices (\$) 0.1638 0.1441 0.1199 0.1607 0.1463 0.1240 0.1838 0.1746 0.0794
Exec. Time (s) 2.469 2.116 1.759 2.356 2.166 1.847 2.812 2.514 1.038
Quality of Plan (QoP) $\uparrow$ 0.1935 0.2224 0.2265 0.2011 0.2201 0.2237 0.1620 0.1929 0.2605
% of Valid Plans $\uparrow$ 100% 100% 100% 100% 100% 100% 100% 100% 100%
Non-Sequential Planning
Task Performance $\uparrow$ 0.2927 0.2630 0.2978 0.4389 0.4408 0.4346 0.4031 0.4270 0.5659
 Costs $\downarrow$ Exec. Prices (\$) 0.1440 0.1231 0.1157 0.2233 0.1628 0.1607 0.1963 0.2011 0.1211
Exec. Time (s) 2.216 1.916 1.447 3.333 2.187 2.204 2.279 2.169 0.7221
Quality of Plan (QoP) $\uparrow$ 0.0193 0.0229 0.0469 0.0225 0.0768 0.0756 0.0284 0.0361 0.1762
% of Valid Plans $\uparrow$ 41.67% 75% 66.67% 100% 100% 100% 100% 100% 100%
Table 2: Compare CATP-LLM with other methods on sequential/non-sequential planning tasks. Arrow $\uparrow$ / $\downarrow$ means higher/lower is better.
@@ 5.1 Setup
Dataset. We use the proposed OpenCATP as the evaluation platform, which features diverse compositional tasks that cannot be solved with a single tool. In particular, we evaluate the efficacy of CATP-LLM in two scenarios: sequential planning and non-sequential planning. For sequential planning, we randomly sample 87 tasks from OpenCATP for evaluation, with 69 for training and 18 for testing. For non-sequential planning, 24 tasks are randomly sampled, with 12 for training and 12 for testing. Note that each task is associated with 40 input data samples of various sizes.
Metrics. Leveraging the OpenCATP platform, we use quality of plan (QoP), task performance scores, and execution prices defined in Eq.(5) as the evaluation metrics. By default, we set the weight parameter in Eq.(5) $\alpha=0.5$. Moreover, we also introduce execution time as the metric to intuitively reflect the overhead of tool plans.
Implementation. When implementing our CATP-LLM framework over the OpenCATP platform, we use the QoP function in Eq.(5) for the reward calculation. That said, the performance scores and costs of the reward function in Eq.(4) are consistent with those of the QoP function in Eq.(5). More details can be found in Appendix §$\S$§C.
Baselines. We compare our CATP-LLM framework with other representative frameworks. Specifically, for prompt engineering based paradigm, we implement the following baselines for comparison: zero-shot learning [19], few-shot learning [28, 40] and HuggingGPT [37]. HuggingGPT is similar to few-shot learning but involves a dedicated planning procedure. It first provides in-context examples to prompt the LLM to decompose a task into sub-tasks in JSON format and next associates each sub-task with a tool for task solving. For the fine-tuning paradigm, the following methods are implemented for comparison: instruction fine-tuning (IFT) [13, 35, 47] and RLTF [10]. IFT refines the LLM’s behavior by fine-tuning it over labeled data containing task descriptions and tool plans. RLTF employs an online RL algorithm to fine-tune the LLM, where the LLM actively devises tool plans and adjusts its behavior based on plan execution feedback. In particular, feedback is replayed to the LLM only after a complete plan is generated. We use QoP in Eq.(5) as the feedback signal.
For the prompt-based methods, we use GPT-3.5 [29] and GPT-4 [1] as the backbone LLMs. For the fine-tuning methods, including CATP-LLM, the open-sourced Llama2-7B [41] is employed as the default LLM. Besides, we use LoRA [14] with rank 64 to fine-tune Llama2-7B for all fine-tuning methods. Furthermore, we adapt our tool planning language for zero-shot, few-shot, IFT and RLTF methods, since we find that they achieve poor performance in non-sequential planning without an appropriate formatted language to support the generation of non-sequential plans.
Hardware. We conduct all experiments on a Intel Xeon Sliver 20C40T server with two 32GB V100 GPUs.
@@ 5.2 Main Results
Sequential Planning. Table 2 compares CATP-LLM with other methods for sequential planning. Both GPT-3.5 and GPT-4 incur high execution costs across various prompt-based learning frameworks, even though they are explicitly prompted with tool cost information to generate tool plans of low costs. In contrast, CATP-LLM effectively reduces the execution costs of tool plans while maintaining comparable performance. Hence, CATP-LLM outperforms both GPT-3.5 and GPT-4 even when using Llama2-7B as its backbone, with 16.5%-34.6% improvements in QoP. Additionally, CATP-LLM also achieves 60.8% and 35.0% higher QoP than IFT and RLTF, respectively. On one hand, IFT fine-tunes the LLM on labeled data of task-plan pairs, which faces difficulty in generalizing to new tasks. On the other hand, unlike CATP-LLM that enables the LLM to evaluate plan costs each time a decision is made, RLTF only allows the LLM to access the plan quality when the plan is completely generated and executed. Consequently, this can mislead the LLM to aggressively optimize plan performance with the price of high execution costs.
Non-sequential Planning. As shown in Table 2, CATP-LLM demonstrates significantly better performance than other methods in non-sequential planning. For instance, compared to GPT-4-powered approaches, CATP-LLM achieves 28.4%-30.2% higher plan performance scores, 24.7%-45.8% lower execution prices, 67.0%-78.3% lower execution time as well as 129.4%-683.0% higher QoP. Additionally, we observe that GPT-3.5 produces many invalid plans (e.g. linking an image-input tool behind a text-output tool) due to the complexity of non-sequential tasks. In comparison, CATP-LLM succeeds to guarantee the validness of tool plans even with the 7B Llama2 as the backbone. It is worth noting that while GPT-3.5 achieves comparable execution prices to CATP-LLM, its execution time is much higher. This is because most plans generated by GPT-3.5 maintain a sequential structure, and thus cannot be executed in parallel to reduce execution time.
Qualitative Results. Figure 1 illustrates the cost-aware tool planning capability of CATP-LLM. For a complex task requiring multiple outputs (i.e. an image caption in German and an object name), CATP-LLM accurately devises a non-sequential plan that can be executed in parallel to reduce execution time. In contrast, the plans generated by GPT-3.5 either follow a sequential structure or lack the necessary machine translation tool, resulting in task failure and low performance scores. While HuggingGPT with GPT-4 includes a costly super-resolution tool, CATP-LLM determines to omit it, as this tool will not significantly enhance performance given that the image size is not very small, but will drastically increase costs. Thus, CATP-LLM efficiently reduces plan costs without sacrificing performance, achieving higher QoP than HuggingGPT.

### v3

@@ 3.2 Cost-Aware Offline Reinforcement Learning
With the proposed TPL, the LLM is capable to devise various tool plans simply by predicting tool and dependency tokens. To optimize the plan generation process, CATP-LLM then fine-tunes the LLM through the efficient CAORL mechanism. Specifically, CAORL includes: i) a context augmentation scheme to effectively integrate tool cost information into the LLM’s input context; ii) an offline RL-based fine-tuning algorithm that guides the LLM to achieve the optimal performance-cost trade-off in tool planning.
@@ 3.2.1 Context Augmentation
Figure 4: Illustrations of the prompt template and context augmentation design. Complete prompts are displayed in Appendix §$\S$§A.
The details of the context augmentation scheme are elaborated as follows. First, we categorize the input sizes into $\{1,\cdots,k\}$ levels as two very close sizes will not make a big difference (e.g., resolution 480$\times$520 vs. 490$\times$510, batch size 32 vs. 36). Subsequently, each tool is assigned with a cost attribute vector $c(t_{i})=(c_{1}^{i},\cdots,c_{k}^{i})$, indicating its execution costs at different levels of input sizes. Afterward, we employ a linear layer to extract features from the tool cost attributes. We then encode the importance of cost features by adding the following importance vector:
@@
The remaining problem here is how to acquire the cost attributes of each tool $c(t_{i})$. To tackle this issue, we manually construct various tool plans and ensure that each tool is used at least once in these plans. We then simulate diverse input data at varying size levels, feed the simulated data to each plan, and record the execution costs of each tool (e.g., execution time and memory consumption), which ultimately yields the cost attributes of each tool.
@@ 3.2.2 Learning Through Offline RL
Despite augmenting the LLM with rich input context, achieving cost-aware tool planning remains challenging due to the inherent trade-off between plan performance and costs. Hence, it is imperative to fine-tune the LLM to acquire the domain knowledge for cost-aware planning. Notably, we observe that the selection of a tool can have cascading effects on the subsequent selection. For instance, selecting a powerful yet costly tool may lead to avoiding the selection of expensive tools later to prevent costs from becoming unaffordable. This suggests that the cost-aware tool planning is essentially a sequential decision problem, a domain where reinforcement learning (RL) [51] excels. Thus, we design the fine-tuning algorithm based on RL, as illustrated in Figure 5 and detailed below.
@@
In lack of public tool plan datasets for fine-tuning, we propose a dataset generation pipeline to collect the necessary training data. We begin by constructing a graph where each node represents a tool and each edge denotes the dependency between two tools. For each task, we then generate a tool plan by searching for a sub-graph within the tool graph that could potentially solve the target task. To reduce the search space, we prune tools that are not relevant to the target task (e.g., image captioning tool for an object detection task). Besides, since not all generated plans are valid (i.e., executable), we will filter out those invalid plans and utilize the rest as the plan dataset for fine-tuning.
4 OpenCATP Dataset
In this section, we elaborate the details of the proposed OpenCATP dataset.
Evaluation tasks. OpenCATP features a range of compositional tasks for the comprehensive evaluation of cost-aware tool planning. It includes 87 sequential tasks that can be solved with sequential plans, which are constructed from the popular sequential planning dataset OpenAGI [15]. These tasks can be categorized as: image-in-image-out tasks (e.g., transforming a blurry image to a regular image) and image-in-text-out tasks (e.g., image classification and captioning). More importantly, it introduces 24 challenging non-sequential tasks, which are absent in previous datasets. These tasks can be classified into two categories: image-in-text-out tasks and image-in-image-text-out tasks (e.g., regularizing a noisy image and identifying the object in it). The detailed construction of sequential/non-sequential tasks is provided in Appendix §$\S$§B, with Table B.1 offering examples of different types of tasks.
Note that each task in OpenCATP is composed of 100 samples of varying input sizes, resulting in 8,700 and 2,400 data points for the evaluation of sequential and non-sequential planning, respectively. Besides, OpenCATP includes 10 open-sourced models as tools (e.g. image captioning model and deblurring model), offering diverse options for tool composition for the LLM to schedule to solve different tasks, as detailed in Appendix §$\S$§B.
Evaluation metrics. Another key feature of OpenCATP is that it defines a cost model to comprehensively reflect the costs of tool plans, inspired by the Function as a Service (FaaS) [44, 26] platforms. To be specific, FaaS providers have developed a mature pricing model to charge for the service of running a function based on the execution time and resource consumption. Hence, we take cues from the FaaS platforms and design a pricing model in OpenCATP to quantify the overall costs of a tool plan, which jointly considers execution time and instant/constant memory consumption (detailed in Appendix §$\S$§B). Based on this model, we introduce a unified metric Quality of Plan (QoP) in OpenCATP to quantify the quality of a tool plan:
@@ 4 OpenCATP Dataset
where $P_{task}(p)$ quantifies the performance of a executed plan on the target task, which can be ViT scores [15] for image-out tasks and BERT scores [63] for text-out tasks; $C_{price}(p)$ calculates the execution prices of a tool plan reflecting its overall costs; $\alpha\in(0,1)$ is the weight parameter. We employ min-max normalization on these two metrics when calculating QoP to make sure their values fall in the same scale. Notably, the min/max values are derived from empirical measurement.
Image-in-Image-out Tasks Image-in-Text-out Tasks Average
Methods Task Scores $\uparrow$ Exec. Prices (\$) $\downarrow$ Exec.   Time (s) $\downarrow$ QoP $\uparrow$ Task Scores $\uparrow$ Exec. Prices (\$) $\downarrow$ Exec.     Time (s) $\downarrow$ QoP $\uparrow$ Task Scores $\uparrow$ Exec. Prices (\$) $\downarrow$ Exec.   Time $\downarrow$ (s) QoP $\uparrow$
Zero-Shot (GPT-3.5) 0.694 0.170 2.457 0.197 0.585 0.134 2.526 0.174 0.639 0.152 2.491 0.186
Few-shot (GPT-3.5) 0.715 0.148 2.073 0.227 0.618 0.123 2.331 0.201 0.666 0.136 2.202 0.214
HuggingGPT (GPT-3.5) 0.674 0.127 1.782 0.226 0.617 0.087 1.656 0.231 0.645 0.107 1.714 0.228
HYDRA (GPT-3.5) 0.558 0.035 0.460 0.248 0.625 0.062 1.144 0.258 0.591 0.049 0.802 0.253
Zero-Shot (GPT-4) 0.699 0.163 2.260 0.206 0.616 0.150 2.844 0.176 0.658 0.156 2.552 0.190
Few-shot (GPT-4) 0.714 0.148 2.098 0.225 0.616 0.133 2.509 0.190 0.665 0.141 2.304 0.208
HuggingGPT (GPT-4) 0.675 0.131 1.893 0.221 0.616 0.085 1.615 0.232 0.646 0.108 1.754 0.227
HYDRA (GPT-4) 0.597 0.050 0.615 0.253 0.623 0.058 1.079 0.259 0.610 0.054 0.847 0.256
IFT (Llama2-7B) 0.616 0.245 4.441 0.092 0.651 0.176 2.619 0.170 0.634 0.210 3.530 0.131
RLTF (Llama2-7B) 0.743 0.234 4.310 0.165 0.687 0.166 2.284 0.196 0.715 0.200 3.297 0.181
CATP-LLM (Llama2-7B) 0.711 0.106 1.866 0.262 0.651 0.074 0.876 0.260 0.681 0.090 1.371 0.261
Table 2: Compare CATP-LLM with other methods on different sequential planning tasks. Arrow $\uparrow$ / $\downarrow$ means higher/lower is better.
Image-in-Text-out Tasks Image-in-Image-Text-out Tasks Average
Methods Task Scores $\uparrow$ Exec. Prices (\$) $\downarrow$ Exec.  Time (s) $\downarrow$ QoP $\uparrow$ Task Scores $\uparrow$ Exec. Prices (\$) $\downarrow$ Exec.  Time (s) $\downarrow$ QoP $\uparrow$ Task Scores $\uparrow$ Exec. Prices (\$) $\downarrow$ Exec.  Time $\downarrow$ (s) QoP $\uparrow$
Zero-Shot (GPT-3.5) 0.506 0.178 2.632 0.095 0.150 0.121 1.938 -0.031 0.328 0.149 2.285 0.032
Few-shot (GPT-3.5) 0.318 0.091 1.686 0.079 0.244 0.133 1.991 0.004 0.281 0.112 1.839 0.041
HuggingGPT (GPT-3.5) 0.349 0.164 2.108 0.029 0.267 0.086 1.049 0.057 0.308 0.125 1.579 0.043
HYDRA (GPT-3.5) 0.184 0.027 0.210 0.068 0.269 0.087 1.387 0.058 0.227 0.057 0.798 0.064
Zero-Shot (GPT-4) 0.654 0.252 3.619 0.105 0.315 0.206 3.169 -0.024 0.485 0.229 3.394 0.040
Few-shot (GPT-4) 0.635 0.197 2.530 0.143 0.330 0.142 1.991 0.039 0.482 0.170 2.261 0.091
HuggingGPT (GPT-4) 0.624 0.199 2.670 0.136 0.326 0.138 1.938 0.040 0.475 0.168 2.304 0.088
HYDRA (GPT-4) 0.326 0.038 0.712 0.129 0.347 0.118 2.005 0.069 0.337 0.077 1.358 0.099
IFT (Llama2-7B) 0.176 0.158 1.485 -0.051 0.516 0.215 2.676 0.068 0.346 0.186 2.080 0.008
RLTF (Llama2-7B) 0.323 0.150 1.400 0.029 0.478 0.226 2.553 0.039 0.401 0.188 1.976 0.034
CATP-LLM (Llama2-7B) 0.641 0.093 0.753 0.238 0.528 0.134 0.706 0.145 0.585 0.114 0.730 0.192
Table 3: Compare CATP-LLM with other methods on different non-sequential planning tasks. Arrow $\uparrow$ / $\downarrow$ means higher/lower is better.
@@ 5.1 Setup
Datasets. For sequential planning, we adopt the sequential tasks of OpenCATP, which are originally constructed from the popular sequential planning dataset OpenAGI [15]. For non-sequential planning, we utilize the tasks developed from OpenCATP.
Metrics. Leveraging the OpenCATP evaluation scheme, we use Quality of Plan (QoP), task performance scores, and execution prices defined in Eq.(5) as the evaluation metrics. By default, we set the weight parameter in Eq.(5) $\alpha=0.5$. Moreover, we also introduce execution time as the metric to intuitively reflect the overhead of tool plans.
Implementation. When implementing our CATP-LLM framework over OpenCATP, we use the QoP function in Eq.(5) for the reward calculation. That said, the performance scores and costs of the reward function in Eq.(4) are consistent with those of the QoP function in Eq.(5). More details can be found in Appendix §$\S$§C.
Baselines. We compare CATP-LLM with other representative frameworks. Specifically, for prompt engineering based paradigm, we implement the following baselines for comparison: zero-shot learning (Zero-Shot) [25], few-shot learning (Few-Shot) [35, 48], HuggingGPT [45] and HYDRA [24]. HuggingGPT is similar to few-shot learning but involves a dedicated planning procedure. It first provides in-context examples to prompt the LLM to decompose a task into sub-tasks in JSON format and next associates each sub-task with a tool for task solving. HYDRA prompts the LLM to generate codes for scheduling tools, where each tool is represented as a python API. For the fine-tuning paradigm, the following methods are implemented for comparison: instruction fine-tuning (IFT) [18, 43, 55] and RLTF [15]. IFT refines the LLM’s behavior by fine-tuning it over labeled data containing task descriptions and tool plans. RLTF employs an online RL algorithm to fine-tune the LLM, where the LLM learns to optimize its decisions based on the feedback of final plan execution. We use QoP in Eq.(5) as the feedback signal for RLTF.
For the prompt-based methods, we use GPT-3.5 [36] and GPT-4 [1] as the backbone LLMs. To ensure fairness, we also provide tool cost information in the prompts for these methods, as detailed in Appendix §$\S$§C. For the fine-tuning methods, including CATP-LLM, the open-sourced Llama2-7B [49] is employed as the default LLM. Besides, we use LoRA [19] with rank 64 to fine-tune Llama2-7B for all fine-tuning methods. Furthermore, we adapt our tool planning language for Zero-Shot, Few-Shot, IFT and RLTF methods, since we find that they achieve poor performance in non-sequential planning without an appropriate formatted language to support the generation of non-sequential plans.
Hardware. We conduct all experiments on an Intel Xeon Sliver 20C40T server with two 32GB V100 GPUs.
@@ 5.2 Main Results
Sequential Planning. Table 2 compares CATP-LLM with other methods for sequential planning. We can see that CATP-LLM consistently maintains low execution costs of tool plans while achieving comparable performance across all cases. Thus, CATP-LLM outperforms GPT-based methods even when using Llama2-7B as its backbone, with 1.5%-40.6% improvements in QoP on average. Additionally, CATP-LLM also achieves 99.0% and 44.3% higher average QoP than IFT and RLTF, respectively. On one hand, IFT fine-tunes the LLM on labeled data of task-plan pairs, which faces difficulty in generalizing to new tasks. On the other hand, unlike CATP-LLM that enables the LLM to evaluate plan costs each time a decision is made, RLTF only allows the LLM to access the plan quality when the plan is completely generated and executed. Consequently, this can mislead the LLM to aggressively optimize plan performance with the price of high execution costs.
Method Task Scores Exec. Prices (\$) % of Valid Plans QoP
w/o context augmentation 0.654 0.092 100% 0.246
w/o intermediate feedback 0.708 0.164 100% 0.209
w/o adaptive masking 0.634 0.086 70.2% 0.242
CATP-LLM 0.681 0.090 100% 0.261
Non-sequential Planning. As shown in Table 3, CATP-LLM demonstrates significantly better performance than other methods in non-sequential planning across all tasks, with 92.0%-375.4% higher QoP on average. Moreover, we empirically observe that GPT-3.5 generates many invalid plans due to the complexity of non-sequential tasks: 41.7% for Zero-Shot, 75% for Few-Shot, and 66.7% for HuggingGPT. For example, it may link an image-input tool behind a text-output tool. In comparison, CATP-LLM succeeds to guarantee the validity of 100% tool plans even with the 7B Llama2 as the backbone. It is worth noting that while some methods (e.g., Zero-Shot and HYDRA) may achieve comparable execution prices to CATP-LLM, their execution time is much higher. This is because most of their plans generated maintain a sequential structure, and thus cannot be executed in parallel to reduce execution time.
Qualitative Results. Figure 1 illustrates the cost-aware tool planning capability of CATP-LLM. For a complex task requiring multiple outputs (i.e. an image caption in German and an object name), CATP-LLM accurately devises a non-sequential plan that can be executed in parallel to reduce execution time. In contrast, the plans generated by GPT-3.5 follow a sequential structure or lack the necessary machine translation tool, resulting in task failure and low performance scores. While HuggingGPT with GPT-4 includes a costly super-resolution tool, CATP-LLM determines to omit it, as this tool will not significantly enhance performance given that the image size is not very small, but will drastically increase costs. Thus, CATP-LLM efficiently reduces plan costs without sacrificing performance, achieving higher QoP than HuggingGPT.